# Quantum-vacuum-induced topological edge polaritons

### v1

@@ V Concluding remarks
In this paper, we uncovered the role of quantum vacuum in generating a pair of topologically protected edge states, termed quantum-vacuum-induced topological edge polaritons, when a single spinful fermion living on a 1D lattice is subject to a TC-like coupling with a single photonic mode. The intuition behind our model construction was first elucidated by considering special parameter values and analytically identifying the exact quantum-vacuum-induced topological edge polaritons. We then demonstrated the robustness of the obtained edge polaritons by performing numerical studies on more general parameter values, as well as by subjecting the system to spatial disorder and counterrotating interaction effect. We further highlighted their quantum vacuum nature by demonstrating that such edge states yield almost zero mean photon number and are otherwise absent in the limit of classical light.
@@
In summary, the quantum-vacuum-induced topological edge polaritons discovered in this paper are made possible due to the fact that the Fock space is bounded from below by the zero photon state, which in turn mimics a physical boundary’s ability of capturing topological effects. It is expected that this simple insight could be adapted to construct a variety of novel topological phases that incorporate Fock space as an artificial dimension. For example, it is imagined that a suitable fermion-photon interaction could be devised to construct the 1D chain analogues of the Chern insulator which supports a unidirectional-moving eigenstate from one edge to the other at very low photon mean number. The possibility and detailed analysis of such a system deserves a separate study that is best left for future work.
@@
Other potential follow-up studies that could benefit from this paper’s findings include considering the effects of non-Hermiticity, few- to many-body fermionic interactions, and/or time-periodicity (additional classical light). In particular, non-Hermitian topological systems form an emerging active research area in recent years that has yielded various interesting phenomena such as the so-called non-Hermitian skin effect Lee16 ; Alvarez18 ; Yao18 ; Kunst18 ; Yao18b ; Li20 and topological effects on the complex energy structure Gong18 ; Lee19 ; Borgnia20 ; Okuma20 ; Zhang20 . As non-Hermiticity serves as an alternative description of open quantum systems, it could be particularly relevant in experiments involving fermions in a relatively low-Q cavity. Meanwhile, the motivation for considering few- to many-body fermionic interaction is twofold. First, many-body interaction effect is known to enrich topological phases, which lead to the emergence of fractional quantum Hall insulators Levin09 ; Neupert11 ; Leonard23 ; Lu24 , Fibonacci anyons Nayak08 ; Lahtinen17 ; Trebst08 ; Mong14 , and parafermions Fendley12 ; Alicea16 ; Clarke13 ; Iemini17 ; Bomantara21 , among others. Second, simultaneous interaction between a single photon mode and multiple fermions is known to yield superradiance, which may have nontrivial effect on the system’s overall topology. Finally, time-periodicity is the main ingredient behind the so-called Floquet topological phases and their unique edge state features that have no static counterparts Bomantara16 ; Rudner13 ; Nathan15 ; Asboth14 ; Ho14 ; Bomantara20 ; Zhou18 ; Zhu22 ; Bomantara22 ; Koor22 . Such edge states could be potentially advantageous for quantum computing applications Bomantara18 ; Bomantara18b ; Bauer19 ; Bomantara20b ; Bomantara20c ; Matthies22 . Given the richness of topological phenomena induced by these effects on their own, combining them with the approach developed here is expected to be even more fruitful for uncovering the full interplay between topology and quantum vacuum.
@@ Acknowledgements.
Appendix A Quantum-vacuum-induced edge polaritons at $E_{Z}\neq 0$ and $J_{1},\omega=0$ up to first order in $J_{2}$
@@
where $f(n)$ is defined in Eq. (27) and $|0_{R}^{(1)}\rangle$ is a correction term that is first-order in $J_{2}$. Using Eqs. (29) and (LABEL:chk4), we then obtain
@@
thus resulting in a term that is first-order in $J_{2}$. In principle, one may continue to appropriately define $|0_{L}^{(1)}\rangle$ and $|0_{R}^{(1)}\rangle$ to obtain both zero modes with higher-order terms in $J_{2}$. However, we choose to stop at first-order approximation to avoid obscuring the main physics with complex expressions.

### v2

@@ V Concluding remarks
In this paper, we uncovered the role of quantum vacuum in generating a pair of topologically protected edge states, termed quantum-vacuum-protected topological edge polaritons, when a single two-level neutral atom or artifical atom living on a 1D lattice is subject to a TC-like coupling with a single photonic mode. The intuition behind our model construction was first elucidated by considering special parameter values and analytically identifying the exact quantum-vacuum-protected topological edge polaritons. We then demonstrated the robustness of the obtained edge polaritons by performing numerical studies on more general parameter values, as well as by subjecting the system to spatial disorder and counterrotating interaction effect. We further highlighted their quantum vacuum nature by demonstrating that such edge states yield almost zero mean photon number and are otherwise absent in the limit of classical light.
@@
In summary, the quantum-vacuum-protected topological edge polaritons discovered in this paper are made possible due to the fact that the Fock space is bounded from below by the zero photon state, which in turn mimics a physical boundary’s ability of capturing topological effects. It is expected that this simple insight could be adapted to construct a variety of novel topological phases that incorporate Fock space as an artificial dimension. For example, it is imagined that a suitable light-matter interaction could be devised to construct the 1D chain analogues of the Chern insulator which supports a unidirectional-moving eigenstate from one edge to the other at very low photon mean number. The possibility and detailed analysis of such systems deserve a separate study that is best left for future work.
Even in the context of the quantum-vacuum-protected edge polaritons discovered here, it is also interesting to explore if they could arise under a different type of light-matter interaction. In particular, the TC-like interaction considered in this paper is typically only relevant for neutral particles, e.g., atoms and excitons. For charged fermions, a more appropriate light-matter interaction is that which arises via Peierls substitution, e.g., in the spirit of Refs. Schlawin22 ; Dmytruk22 ; Bacciconi24 . Such an interaction is not expected to yield quantum-vacuum-protected edge polaritons when directly applied to the particle Hamiltonian considered here, since it only leads to terms that couple either (pseudo)spin and photonic degrees of freedom or sublattice and photonic degrees of freedom as opposed to the interaction term considered in this paper that couples all pseudospin, sublattice, and photonic degrees of freedom simultaneously. Nevertheless, applying Peierls substitution on a different fermionic lattice model could potentially generate the necessary fermion-photon interaction that supports quantum-vacuum-protected edge polaritons and may in fact serve as a good alternative to simulate our system without RWA. To this end, the analogy with the second-order topological insulator as pointed out in this paper could be a key idea towards identifying the right fermionic lattice model, which would then constitute a promising follow-up study.
@@
Other potential future directions that could benefit from this paper’s findings include considering the effects of non-Hermiticity, few- to many-body particle interactions, and/or time-periodicity (additional classical light). In particular, non-Hermitian topological systems form an emerging active research area in recent years that has yielded various interesting phenomena such as the so-called non-Hermitian skin effect Lee16 ; Alvarez18 ; Yao18 ; Kunst18 ; Yao18b ; Li20 and topological effects on the complex energy structure Gong18 ; Lee19 ; Borgnia20 ; Okuma20 ; Zhang20 . As non-Hermiticity serves as an alternative description of open quantum systems, it could be particularly relevant in experiments involving particles in a relatively low-Q cavity. Meanwhile, the motivation for considering few- to many-body particle interaction is twofold. First, many-body interaction effect is known to enrich topological phases, which leads to the emergence of fractional quantum Hall insulators Levin09 ; Neupert11 ; Leonard23 ; Lu24 , Fibonacci anyons Nayak08 ; Lahtinen17 ; Trebst08 ; Mong14 , and parafermions Fendley12 ; Alicea16 ; Clarke13 ; Iemini17 ; Bomantara21 , among others. Second, simultaneous interaction between a single photon mode and multiple particles is known to yield superradiance, which may have nontrivial effect on the system’s overall topology. Finally, time-periodicity is the main ingredient behind the so-called Floquet topological phases and their unique edge state features that have no static counterparts Bomantara16 ; Rudner13 ; Nathan15 ; Asboth14 ; Ho14 ; Bomantara20 ; Zhou18 ; Zhu22 ; Bomantara22 ; Koor22 . Such edge states are further known to be potentially advantageous for quantum computing applications Bomantara18 ; Bomantara18b ; Bauer19 ; Bomantara20b ; Bomantara20c ; Matthies22 . Given the richness of topological phenomena induced by these effects on their own, combining them with the approach developed here is expected to be even more fruitful for uncovering the full interplay between topology and quantum vacuum.
@@ Acknowledgements.
Appendix A Quantum-vacuum-protected edge polaritons at $E_{Z}\neq 0$ and $J_{1},\omega=0$ up to first order in $J_{2}$
@@
where $f(n)$ is defined in Eq. (27) and $|0_{R}^{(1)}\rangle$ is a correction term that is first-order in $J_{2}$. Using Eqs. (29) and (30), we then obtain
@@
thus resulting in a term that is first-order in $J_{2}$. In principle, one may continue to appropriately define $|0_{L}^{(1)}\rangle$ and $|0_{R}^{(1)}\rangle$ to obtain both zero modes with higher-order terms in $J_{2}$. However, we choose to stop at the first-order approximation to avoid obscuring the main physics with complex expressions.